# MicroEvoEval: A Systematic Evaluation Framework for Image-Based Microstructure Evolution Prediction

### v1

@@ 3.1 The Taxonomy of Microstructure Evolution Tasks
In this work, we select representative applications for each of these PDE categories. Specifically, we consider plane-wave propagation for periodic structures, grain growth governed by the A-C equation, spinodal decomposition described by the C–H equation, and dendritic solidification as a coupled multi-physics process. It is worth noting that generating high-quality data for these PDEs is nontrivial. To ensure reproducibility and enable meaningful benchmark comparisons, we adopt the datasets and PDE formulations introduced in (yang2021self), which are widely used in the modeling of microstructural evolution (Fan2024; jing2025research). We then applied specific processing to these datasets to align them with our evaluation goals in the next subsection. In the following, we first briefly describe each task and present its associated governing equation. [^1] More details of the physical background and derivation of these tasks are shown in Technical Appendix A.
@@ 3.2 Tasks and Datasets
To investigate the relationship between short-term accuracy and long-term stability, we design a dual evaluation setting based on our curated datasets. The short-term setting assesses a model’s ability to capture immediate dynamics (e.g., predicting the next 10 frames from 10 inputs), while the long-term setting evaluates autoregressive performance over extended horizons (up to 50 and 90 future frames). [^2] A key area of focus in this domain is assessing the performance of models when they are trained for short-term prediction but then deployed autoregressively for long-term forecasting. This twofold structure enables a systematic analysis of error accumulation and the extent to which short-term accuracy correlates with long-term reliability. Figure 1 summarizes the benchmark tasks, datasets, and key statistics in MicroEvoEval.  [^3] More implementation details of our datasets are shown in Code and Data Appendix D.1.
@@
For this task, we employ datasets from simulations of dendritic solidification based on Equation (6). The original study provides 940 training simulations, each producing 100 frames. Our training set is built from overlapping clips extracted from these simulations, resulting in 19740 samples. The provided test simulations, which notably include out-of-distribution samples to assess generalization, are partitioned by us into evaluation sets for short-term (2100 test cases) and long-term (1100 test cases for predicting next 50 frames and 100 test cases for predicting next next 90 frames) performance assessment.
@@ 3.3 Metrics
To enable a holistic assessment, our evaluation framework integrates metrics across three dimensions: predictive accuracy, physical fidelity, and computational efficiency, each critical for practical applications in materials science. [^4] More implementation details of our metrics are shown in Code and Data Appendix D.2.
@@
For tasks where the overall phase fraction is a key evolving property, such as plane-wave propagation, spinodal decomposition, and dendritic solidification, we introduce the Log-Error of Total Area Proportion (L-ETAP). [^5] To better compare performance across different orders of magnitude, as errors can vary significantly between models, we report these metrics on a base-10 logarithmic scale. This metric is physically significant as it evaluates a model’s ability to conserve the total mass or volume of the evolving phase, a fundamental physical constraint. Specifically, we track this by computing the phase area fraction $A(\Omega_{i,t})/S$, where $A(\Omega_{i,t})$ is the total area of the evolving phase (e.g., the white pixels for dendrites) and $S$ is the total image area. Any deviation in this fraction indicates unphysical mass gain or loss over time. L-ETAP therefore captures the degree to which a model maintains global conservation laws across a temporal sequence. It is defined as:
@@ 4.1 Experimental Setup
To conduct this evaluation, we assess a diverse set of 14 models, comprising 5 architectures specifically tailored for MicroEvo and 9 state-of-the-art general-purpose spatio-temporal models. The domain-specific MicroEvo models include E3D-LSTM (yang2021self), ConvGRU (lanzoni2022morphological), PredRNN (farizhandi2023spatiotemporal), ConvLSTM (mao2024spatiotemporal), and VMamba (jing2025research). The general-purpose spatio-temporal architectures consist of PredRNN++ (Wang2018predrnn++), MAU (chang2021mau), SimVP (gao2022simvp), PredRNNv2 (wang2022predrnn), TAU (tan2023temporal), SwinLSTM (tang2023swinlstm), VMRNN (tang2024vmrnn), SimVP.v2 (tan2025simvpv2), and PredFormer (tang2024predformer). All models were trained for a maximum of 200, 300, 300, and 400 epochs of each task with the best performing checkpoint selected based on performance in a held-out validation set. [^7] Further implementation details of each model can be found in Code and Data Appendix E.
@@ 4.2 Main Results
Our experiments provide a deep understanding of MicroEvo prediction by investigating four core questions: (1) whether a model’s short-term accuracy is a reliable indicator of its long-term stability; (2) whether standard image metrics (numerical accuracy) are sufficient for evaluation compared to our curated metrics for physical fidelity; (3) how domain-specific MicroEvo models perform relative to general-purpose spatiotemporal architectures; and (4) which architectural classes offer the best trade-off between performance and computational efficiency. [^8] For more detailed results with additional numerical accuracy metrics, please refer to Technical Appendix B.1.
@@ 4.3 Visual Analysis
In addition to quantitative metrics, we use visual analysis to explore long-term stability and to show the findings of our physical fidelity metrics. [^9] More cases are shown in Technical Appendix B.2.
@@ Appendix A More Details of the Physical Background
Appendix B More Results
B.1 More Details of the Main Results
@@
Model domain Model MAE MSE RMSE SSIM L-ETAP MAE MSE RMSE SSIM L-EASPR MAE MSE RMSE SSIM L-ETAP MAE MSE RMSE SSIM L-ETAP
@@
MAE and MSE Metrics Further Confirm the Short- vs. Long-Term Gap. The inclusion of MAE and MSE in Table 4 reinforces the key finding that short-term performance is a poor indicator of long-term stability. As seen in the table, both MAE and MSE values increase by one to two orders of magnitude for nearly all models when shifting from the 10-frame to the 90-frame prediction horizon. Furthermore, we observe subtle divergences between MAE and MSE rankings. For instance, in the 90-frame prediction for Spinodal Decomposition, TAU achieves the best MAE (0.00310), while VMamba secures the best MSE (0.00047). This is because MAE (L1 norm) is less sensitive to large, outlier pixel errors than MSE (L2 norm).

### v2

@@ Abstract
Codehttps://github.com/ArcueidCroft/MircoEvoEval
Datasetshttps://huggingface.co/datasets/ArcueidCroft/MicroEvoEval
@@ 3.1 The Taxonomy of Microstructure Evolution Tasks
In this work, we select representative applications for each of these PDE categories. To ensure systematic coverage, the selected four tasks follow a physical taxonomy encompassing periodic, non-conserved, conserved, and coupled multi-physics mechanisms. Each task was designed to be non-overlapping in its governing dynamics and data domain.  [^1] We adopt PDE-based datasets for their scalability and clean ground truths, as real-world microstructure evolution data remain scarce and noisy for standardized benchmarking. Specifically, we consider plane-wave propagation for periodic structures, grain growth governed by the A-C equation, spinodal decomposition described by the C–H equation, and dendritic solidification as a coupled multi-physics process. It is worth noting that generating high-quality data for these PDEs is nontrivial. To ensure reproducibility and enable meaningful benchmark comparisons, we adopt the datasets and PDE formulations introduced in (yang2021self), which are widely used in the modeling of microstructural evolution (Fan2024; jing2025research). We then applied specific processing to these datasets to align them with our evaluation goals in the next subsection. In the following, we first briefly describe each task and present its associated governing equation. [^2] More details of the physical background and derivation of these tasks are shown in Technical Appendix A. We also constructed a Dumbbell dataset as extended version.
@@ 3.2 Tasks and Datasets
To investigate the relationship between short-term accuracy and long-term stability, we design a dual evaluation setting based on our curated datasets. The short-term setting assesses a model’s ability to capture immediate dynamics (e.g., predicting the next 10 frames from 10 inputs), while the long-term setting evaluates autoregressive performance over extended horizons (up to 50 and 90 future frames). [^3] A key area of focus in this domain is assessing the performance of models when they are trained for short-term prediction but then deployed autoregressively for long-term forecasting. More implementation details are shown in Code and Data Appendix D.1. This twofold structure enables a systematic analysis of error accumulation and the extent to which short-term accuracy correlates with long-term reliability. Figure 1 summarizes the benchmark tasks, datasets, and key statistics in MicroEvoEval.
@@
For this task, we employ datasets from simulations of dendritic solidification based on Equation (6). The original study provides 940 training simulations, each producing 100 frames. Our training set is built from overlapping clips extracted from these simulations, resulting in 19740 samples. The provided test simulations, which notably include out-of-distribution samples to assess generalization, are partitioned by us into evaluation sets for short-term (2100 test cases) and long-term (1100 test cases for predicting next 50 frames and 100 test cases for predicting next 90 frames) performance assessment.
@@ 3.3 Metrics
To enable a holistic assessment, our evaluation framework integrates metrics across three dimensions: predictive accuracy, physical fidelity, and computational efficiency, each critical for practical applications in materials science. [^4] More implementation details are shown in Code and Data Appendix D.2.
@@
For tasks where the overall phase fraction is a key evolving property, such as plane-wave propagation, spinodal decomposition, and dendritic solidification, we introduce the Log-Error of Total Area Proportion (L-ETAP). This metric is physically significant as it evaluates a model’s ability to conserve the total mass or volume of the evolving phase, a fundamental physical constraint. Specifically, we track this by computing the phase area fraction $A(\Omega_{i,t})/S$, where $A(\Omega_{i,t})$ is the total area of the evolving phase (e.g., the white pixels for dendrites) and $S$ is the total image area. Any deviation in this fraction indicates unphysical mass gain or loss over time. L-ETAP therefore captures the degree to which a model maintains global conservation laws across a temporal sequence. It is defined as:
@@ 4.1 Experimental Setup
To conduct this evaluation, we assess a diverse set of 14 models, comprising 5 architectures specifically tailored for MicroEvo and 9 state-of-the-art general-purpose spatio-temporal models. The domain-specific MicroEvo models include E3D-LSTM (yang2021self), ConvGRU (lanzoni2022morphological), PredRNN (farizhandi2023spatiotemporal), ConvLSTM (mao2024spatiotemporal), and VMamba (jing2025research). The general-purpose spatio-temporal architectures consist of PredRNN++ (Wang2018predrnn++), MAU (chang2021mau), SimVP (gao2022simvp), PredRNN.v2 (wang2022predrnn), TAU (tan2023temporal), SwinLSTM (tang2023swinlstm), VMRNN (tang2024vmrnn), SimVP.v2 (tan2025simvpv2), and PredFormer (tang2024predformer). All models were trained for a maximum of 200, 300, 300, and 400 epochs of each task with the best performing checkpoint selected based on performance in a held-out validation set. [^6] Further implementation details of each model can be found in Code and Data Appendix E.
@@ 4.2 Main Results
Our experiments provide a deep understanding of MicroEvo prediction by investigating four core questions: (1) whether a model’s short-term accuracy is a reliable indicator of its long-term stability; (2) whether standard image metrics (numerical accuracy) are sufficient for evaluation compared to our curated metrics for physical fidelity; (3) how domain-specific MicroEvo models perform relative to general-purpose spatiotemporal architectures; and (4) which architectural classes offer the best trade-off between performance and computational efficiency. [^7] Additional accuracy metrics, performance-cost figure, statistical variance analyses, cases and out-of-distribution evaluations are provided in Technical Appendix B.
@@ 4.3 Visual Analysis
In addition to quantitative metrics, we use visual analysis to explore long-term stability and to show the findings of our physical fidelity metrics.
@@ 5 Conclusion
Acknowledgements
This research is supported by the National Natural Science Foundation of China (No. 11971336, 12571435).
@@ Appendix A More Details of the Physical Background
As part of our ongoing commitment to expanding the benchmark for a more holistic evaluation, we are now introducing an additional dataset.
Dumbbell Evolution via the Allen–Cahn Phase-Field Model
The additional benchmark examines curvature-driven geometric evolution using the Allen–Cahn phase-field formulation. The system models the interface motion under mean curvature flow, where the normal velocity equals the local geometric curvature (deckelnick2005computation). A smooth order parameter $u(\mathbf{x},t)\in[0,1]$ distinguishes the two phases, and the interfacial energy is represented by the diffuse-interface functional
$P_{\varepsilon}(u)=\int_{\mathbb{R}^{d}}\left[\varepsilon\frac{|\nabla u|^{2}}{2}+\frac{1}{\varepsilon}W(u)\right]\,d\mathbf{x},$ (23)
where $\varepsilon>0$ controls the interfacial thickness and $W(u)=\tfrac{1}{2}u^{2}(1-u)^{2}$ is a symmetric double-well potential.
Taking the $L^{2}$–gradient flow of $P_{\varepsilon}$ yields the classical Allen–Cahn equation
$\frac{\partial u}{\partial t}=\Delta u-\frac{1}{\varepsilon^{2}}W^{\prime}(u),$ (24)
which drives the diffuse interface to move with velocity proportional to its mean curvature. In the sharp-interface limit $\varepsilon\to 0$, the zero level set of $u$ evolves according to mean curvature flow, with normal velocity $V_{n}=H$.
For this Dumbbell task, the initial condition is generated from the signed distance function to a two-lobed dumbbell-shaped domain $\Omega(0)$:
$u_{\varepsilon}(\mathbf{x},0)=q\!\left(\frac{d(\mathbf{x},\Omega(0))}{\varepsilon}\right),~~q(s)=\tfrac{1}{2}\left(1-\tanh\!\left(\frac{s}{2}\right)\right),$ (25)
where $d(\mathbf{x},\Omega(0))$ denotes the signed distance. The function $q$ represents the optimal 1D transition profile minimizing the interfacial energy (bretin2022learning).
During evolution, curvature-driven motion causes the narrow neck connecting the two lobes to progressively thin and ultimately pinch off, producing two separate droplets. This process captures geometric singularity formation (topological transition) and curvature-induced interface dynamics within the phase-field paradigm, making it a compelling extended example for our benchmark.
@@
Model domain Model MAE MSE RMSE SSIM L-ETAP MAE MSE RMSE SSIM L-EAPSR MAE MSE RMSE SSIM L-ETAP MAE MSE RMSE SSIM L-ETAP
@@
Table 5: The Variance of the Main Results.
Model domain Model Var(MAE) Var(MSE) Var(RMSE) Var(SSIM) Var(MAE) Var(MSE) Var(RMSE) Var(SSIM) Var(MAE) Var(MSE) Var(RMSE) Var(SSIM) Var(MAE) Var(MSE) Var(RMSE) Var(SSIM)
Short term prediction (10-10)
Microstructural E3DLSTM 0.000082 0.000001 0.000066 0.000077 0.000013 0.000001 0.000087 0.000021 0.000000 0.000000 0.000013 0.000000 0.000000 0.000000 0.000003 0.000000
ConvGRU 0.000014 0.000000 0.000021 0.000001 0.000002 0.000000 0.000044 0.000004 0.000000 0.000000 0.000008 0.000000 0.000000 0.000000 0.000001 0.000000
PredRNN 0.000002 0.000000 0.000004 0.000000 0.000014 0.000001 0.000092 0.000013 0.000000 0.000000 0.000007 0.000000 0.000001 0.000000 0.000008 0.000000
ConvLSTM 0.000032 0.000000 0.000023 0.000000 0.000017 0.000001 0.000091 0.000021 0.000000 0.000000 0.000013 0.000000 0.000000 0.000000 0.000001 0.000000
VMamba 0.000004 0.000000 0.000006 0.000000 0.000001 0.000000 0.000026 0.000002 0.000000 0.000000 0.000011 0.000000 0.000000 0.000000 0.000001 0.000000
Spatiotemporal PredRNN++ 0.000003 0.000000 0.000004 0.000000 0.000015 0.000001 0.000088 0.000019 0.000000 0.000000 0.000009 0.000000 0.000000 0.000000 0.000002 0.000000
MAU 0.000040 0.000001 0.000089 0.000132 0.000024 0.000002 0.000083 0.000055 0.000000 0.000000 0.000033 0.000003 0.000000 0.000000 0.000002 0.000000
SimVP 0.000060 0.000001 0.000086 0.000091 0.000010 0.000001 0.000077 0.000001 0.000000 0.000000 0.000014 0.000001 0.000000 0.000000 0.000002 0.000000
PredRNN.v2 0.000001 0.000000 0.000002 0.000000 0.000023 0.000001 0.000111 0.000038 0.000000 0.000000 0.000015 0.000001 0.000000 0.000000 0.000001 0.000000
TAU 0.000012 0.000000 0.000016 0.000002 0.000005 0.000000 0.000080 0.000009 0.000000 0.000000 0.000019 0.000001 0.000000 0.000000 0.000002 0.000000
SwinLSTM 0.000001 0.000000 0.000022 0.000009 0.000022 0.000001 0.000122 0.000053 0.000001 0.000000 0.000017 0.000003 0.000000 0.000000 0.000002 0.000000
VMRNN 0.000000 0.000000 0.000000 0.000000 0.000008 0.000000 0.000062 0.000010 0.000000 0.000000 0.000008 0.000000 0.000000 0.000000 0.000001 0.000000
SimVP.v2 0.000047 0.000000 0.000062 0.000023 0.000008 0.000000 0.000099 0.000013 0.000000 0.000000 0.000012 0.000001 0.000001 0.000000 0.000007 0.000000
PredFormer 0.000049 0.000000 0.000068 0.000089 0.000014 0.000001 0.000115 0.000036 0.000001 0.000000 0.000027 0.000015 0.000000 0.000000 0.000013 0.000008
Long term prediction (10-50)
Microstructural E3DLSTM 0.000330 0.000016 0.000688 0.002110 0.000060 0.000014 0.000207 0.001009 0.000002 0.000001 0.000145 0.000019 0.000005 0.000002 0.000156 0.000190
ConvGRU 0.000316 0.000005 0.000519 0.000818 0.000028 0.000007 0.000165 0.000444 0.000003 0.000001 0.000162 0.000026 0.000005 0.000001 0.000179 0.000120
PredRNN 0.000062 0.000000 0.000133 0.000120 0.000028 0.000017 0.000068 0.000144 0.000002 0.000001 0.000123 0.000017 0.000200 0.000133 0.000456 0.000476
ConvLSTM 0.000993 0.000066 0.001356 0.001690 0.000110 0.000021 0.000349 0.001380 0.000020 0.000002 0.000023 0.000025 0.000011 0.000003 0.000319 0.000303
VMamba 0.000074 0.000000 0.000125 0.000149 0.000018 0.000003 0.000357 0.000295 0.000002 0.000001 0.000107 0.000019 0.000001 0.000000 0.000038 0.000006
Spatiotemporal PredRNN++ 0.000055 0.000000 0.000110 0.000124 0.000102 0.000019 0.000333 0.001253 0.000004 0.000001 0.000134 0.000031 0.000003 0.000000 0.000100 0.000050
MAU 0.000412 0.000019 0.000912 0.001195 0.000075 0.000020 0.000137 0.001513 0.000013 0.000005 0.000235 0.000162 0.000028 0.000015 0.000454 0.000903
SimVP 0.000689 0.000045 0.001064 0.005579 0.000085 0.000015 0.000312 0.001191 0.000009 0.000001 0.000099 0.000020 0.000006 0.000001 0.000227 0.000153
PredRNN.v2 0.000028 0.000000 0.000057 0.000077 0.000055 0.000014 0.000294 0.000986 0.000005 0.000002 0.000214 0.000049 0.000018 0.000007 0.000483 0.000535
TAU 0.000191 0.000004 0.000308 0.000808 0.000071 0.000014 0.000333 0.000977 0.000004 0.000001 0.000208 0.000036 0.000003 0.000000 0.000119 0.000067
SwinLSTM 0.000399 0.000015 0.000614 0.003368 0.000141 0.000029 0.000317 0.001997 0.000003 0.000001 0.000199 0.000069 0.000010 0.000002 0.000401 0.000228
VRMNN 0.000042 0.000000 0.000098 0.000114 0.000052 0.000009 0.000357 0.000700 0.000004 0.000002 0.000187 0.000043 0.000005 0.000001 0.000217 0.000086
SimVP.v2 0.000458 0.000020 0.000702 0.002441 0.000085 0.000016 0.000341 0.001058 0.000004 0.000001 0.000185 0.000031 0.000007 0.000000 0.000151 0.000040
PredFormer 0.000392 0.000011 0.000571 0.001709 0.000106 0.000020 0.000448 0.001678 0.000016 0.000005 0.000258 0.000169 0.000031 0.000013 0.000753 0.000813
Long term prediction (10-90)
Microstructural E3DLSTM 0.000993 0.000113 0.000121 0.017460 0.000077 0.000022 0.000197 0.002259 0.000012 0.000005 0.000391 0.000135 0.000175 0.000098 0.000888 0.001525
ConvGRU 0.000934 0.000045 0.001557 0.010206 0.000054 0.000016 0.000191 0.001389 0.000011 0.000005 0.000357 0.000098 0.000072 0.000026 0.000798 0.001251
PredRNN 0.000224 0.000004 0.000422 0.002296 0.000045 0.000050 0.000097 0.000041 0.000010 0.000004 0.000291 0.000105 0.000348 0.000318 0.000515 0.001471
ConvLSTM 0.001065 0.000096 0.002163 0.021800 0.000158 0.000038 0.000384 0.003488 0.000022 0.000009 0.000384 0.000158 0.000257 0.000127 0.001931 0.004031
VMamba 0.000221 0.000003 0.000366 0.001609 0.000062 0.000014 0.000816 0.001460 0.000008 0.000003 0.000241 0.000074 0.000018 0.000004 0.000322 0.000219
Spatiotemporal PredRNN++ 0.000341 0.000009 0.000556 0.001898 0.000155 0.000037 0.000392 0.003358 0.000015 0.000004 0.000282 0.000144 0.000035 0.000013 0.000427 0.000653
MAU 0.001692 0.000126 0.001923 0.013549 0.000069 0.000023 0.000110 0.001765 0.000045 0.000022 0.000497 0.000726 0.000134 0.000106 0.000680 0.001725
SimVP 0.001288 0.000088 0.001951 0.019977 0.000018 0.000041 0.000432 0.003019 0.000011 0.000004 0.000331 0.000076 0.000076 0.000023 0.001056 0.001282
PredRNN.v2 0.000139 0.000004 0.000276 0.002563 0.000099 0.000023 0.000392 0.002484 0.000016 0.000010 0.000308 0.000103 0.000282 0.000155 0.001704 0.003899
TAU 0.000483 0.000019 0.000751 0.005430 0.000111 0.000028 0.000355 0.002671 0.000012 0.000005 0.000417 0.000116 0.000038 0.000009 0.000587 0.000558
SwinLSTM 0.000934 0.000066 0.001661 0.010421 0.000157 0.000038 0.000249 0.002839 0.000009 0.000004 0.000233 0.000446 0.000166 0.000067 0.002043 0.002812
VRMNN 0.000409 0.000018 0.000901 0.005297 0.000104 0.000025 0.000551 0.002479 0.000015 0.000006 0.000393 0.000152 0.000095 0.000035 0.001263 0.001480
SimVP.v2 0.000947 0.000063 0.001427 0.011695 0.000123 0.000030 0.000338 0.002850 0.000012 0.000005 0.000397 0.000110 0.000051 0.000012 0.000764 0.000810
PredFormer 0.000979 0.000072 0.001565 0.011022 0.000145 0.000036 0.000467 0.003466 0.000048 0.000023 0.000488 0.000522 0.000251 0.000136 0.002075 0.004706
Figure 4: The performance-cost of Plane-wave propagation.
Figure 5: The performance-cost of Grain growth.
Figure 6: The performance-cost of Spinodal decomposition
Figure 7: The performance-cost of Dendritic Solidification
Appendix B More Results
B.1 More Details of the Main Results
@@
MAE and MSE Metrics Further Confirm the Short- vs. Long-Term Gap. The inclusion of MAE and MSE in Table 4 (with variance, Table 5) reinforces the key finding that short-term performance is a poor indicator of long-term stability. As seen in the table, both MAE and MSE values increase by one to two orders of magnitude for nearly all models when shifting from the 10-frame to the 90-frame prediction horizon. Furthermore, we observe subtle divergences between MAE and MSE rankings. For instance, in the 90-frame prediction for Spinodal Decomposition, TAU achieves the best MAE (0.00310), while VMamba secures the best MSE (0.00047). This is because MAE (L1 norm) is less sensitive to large, outlier pixel errors than MSE (L2 norm).
@@
Except these results from the Table, we also use Figure 4-7 to show the relationship of models between performance and cost, which further confirm the best results of VMamba.
@@ B.2 More Cases
B.3 Main Results of Dumbbell Dataset
To evaluate performance on our new task, we selected representative models from both the microstructural category (PredRNN, VMamba) and the spatiotemporal category (SimVP.v2). The detailed configuration of the dataset can be found in Appendix D.3. Due to the unique nature of the dumbbell evolution, which primarily focuses on the speed of motion rather than phase fraction change or grain coarsening, the existing physical fidelity metrics are not directly applicable. Therefore, our analysis for this task focuses on numerical accuracy metrics.
The Short- vs. Long-Term Performance Gap Persists. Consistent with our primary results, the Dumbbell dataset reaffirms the significant gap between short- and long-term prediction accuracy. For instance, as shown in Table 6, the RMSE for the PredRNN model degrades from 0.000415 in the short-term (10-10) prediction to 0.030105 in the long-term (10-90) prediction. This trend holds for all tested models, providing external confirmation that short-term accuracy is an unreliable indicator of a model’s long-term stability.
Nuanced Performance of VMamba Reveals Task-Specific Strengths. Interestingly, this dataset highlights a scenario where the otherwise dominant VMamba model shows a less pronounced advantage. While it still performs well, it is consistently outperformed by SimVP.v2 in terms of numerical accuracy across all prediction horizons. We hypothesize that this is because the dumbbell evolution involves visually simpler geometric structures compared to tasks like grain growth. The primary challenge here is not modeling complex topological changes, but rather accurately capturing the velocity of the interface evolution. In such a scenario, the sophisticated architectural priors of VMamba may not offer a distinct advantage over a highly efficient, convolution-based model like SimVP.v2, which excels at learning more direct pixel-level transformations. This suggests that even state-of-the-art architectures have specific conditions under which they are most effective.
Table 6: Main Results of Dumbbell Dataset
Dumbbell
Model domain Model MAE MSE RMSE SSIM
Short term prediction (10-10)
Microstructural PredRNN 0.000087 0.000000 0.000415 0.999993
VMamba 0.000346 0.000000 0.000548 0.999973
Spatiotemporal SimVP.v2 0.000063 0.000000 0.000229 0.999997
Long term prediction (10-50)
Microstructural PredRNN 0.000920 0.000116 0.010777 0.997161
VMamba 0.000469 0.000004 0.001897 0.999776
Spatiotemporal SimVP.v2 0.000100 0.000000 0.000679 0.999977
Long term prediction (10-90)
Microstructural PredRNN 0.002772 0.000906 0.030105 0.986203
VMamba 0.000725 0.000042 0.006488 0.998383
Spatiotemporal SimVP.v2 0.000198 0.000006 0.002472 0.999717
B.4 Results of Out of Distribution Experiment
To assess the generalization capabilities of the models beyond the training distribution, we conducted an out-of-distribution (OOD) experiment. This involved testing the pre-trained models on a version of the Dumbbell dataset with physical parameters shifted outside the range seen during training, as detailed in Appendix D.3.
The results highlight that performance degradation under OOD conditions is a persistent challenge. As shown in Table 7, all models experienced an increase in numerical error when faced with this distributional shift. For instance, while SimVP.v2 remains the top-performing model, its long-term (10-50) RMSE nearly doubles from 0.000679 on the in-distribution test set (Table 6) to 0.001379 on the OOD test set (Table 7). This performance drop, even on a task with relatively simple dynamics, underscores that out-of-distribution generalization is a fundamental challenge for current data-driven models. Therefore, developing models with improved physical grounding and resilience to distributional shifts remains a critical and open direction for future research in data-driven materials science.
Table 7: OOD Results of Dumbbell Dataset
Dumbbell
Model domain Model MAE MSE RMSE SSIM
Short term prediction (10-10)
Microstructural PredRNN 0.000118 0.000002 0.001350 0.999987
VMamba 0.000448 0.000001 0.001140 0.999943
Spatiotemporal SimVP.v2 0.000090 0.000000 0.000548 0.999994
Long term prediction (10-50)
Microstructural PredRNN 0.000779 0.000096 0.009812 0.997378
VMamba 0.000552 0.000006 0.002449 0.999696
Spatiotemporal SimVP.v2 0.000156 0.000002 0.001379 0.999915
Long term prediction (10-90)
Microstructural PredRNN 0.002061 0.000572 0.023913 0.990124
VMamba 0.000725 0.000042 0.006488 0.998383
Spatiotemporal SimVP.v2 0.000265 0.000008 0.002782 0.999598